# A variational approach to dimension-free self-normalized concentration

### v1

@@ 7 Summary
The story does not end here. For one, the optimality (or suboptimality) of our bounds has not been established. Can we remove the dependence on $H_{c,\tau}$ in Theorem 4.3, for instance? Is the function $g_{\psi}$ in Theorem 4.1 necessary, or is it simply an artifact of the proof technique? While $g_{\psi}$ is easy to compute, its presence makes it tougher to reason about these bounds analytically.
@@ Appendix A Method of Mixtures for sub-Gaussian Processes
Since $\int N_{t}(\theta)\textnormal{d}\nu(\theta)$ is a supermartingale by Fubini’s theorem, $M_{t}$ remains upper bounded by a nonnegative supermartingale. We may thus apply Ville’s inequality to obtain $P(M_{\tau}\geq 1/\delta)\leq\mathbb{E}[M_{1}]\delta\leq\delta$. In other words, with probability $1-\delta$, $\log M_{\tau}\leq 1/\delta$, which translates to
@@ Appendix B Omitted Proofs
$\int\log Z_{\tau}(\theta)\rho(\textnormal{d}\theta)$ $=\lambda\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}^{2}-\psi_{N}(\lambda)\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}V_{\tau}\widehat{V}_{\tau}^{-1}}^{2}-\psi_{N}(\lambda)\text{Tr}(V_{\tau}\Sigma_{\rho})$
@@
$=\frac{1}{2}\left\{\text{Tr}(\Sigma_{\nu}^{-1}\Sigma_{\rho})+\|S_{\tau}\|^{2}_{\widehat{V}_{\tau}^{-1}\Sigma_{\nu}^{-1}\widehat{V}_{\tau}^{-1}}-d+\log\left(\frac{\det\Sigma_{\nu}}{\det\Sigma_{\rho}}\right)\right\}$
@@
Let $(S_{t},V_{t})$ be a sub-$\psi$ process and set
$Z_{t}(\theta)=\exp\left\{\lambda\langle\theta,S_{t}\rangle-\psi(\lambda)\langle\theta,V_{t}\theta\rangle\right\}.$ (B.5)
For any $\theta\in\mathbb{S}^{d-1}$, the process $(Z_{t}(\theta))$ is upper bounded by some nonnegative supermartingale by assumption. We claim that the same holds for all $\theta\in\mathbb{B}^{d}$. This is shown in Chugg et al. [14, Lemma 2.10], but let us prove it for completeness. For any $\theta\in\mathbb{B}^{d}$, let $\phi=\theta/\|\theta\|$ and notice that
$\exp\{\lambda\langle\theta,S_{t}\rangle-\psi(\lambda)\langle\theta,V_{t}\theta\rangle\}$
$=\exp\{\lambda\|\theta\|\langle\phi,S_{t}\rangle-\psi(\lambda)\|\theta\|^{2}\langle\phi,V_{t}\phi\rangle\}$
$\leq\exp\{\lambda\|\theta\|\langle\phi,S_{t}\rangle-\phi(\|\theta\|\lambda)\langle\phi,V_{t}\phi\rangle\},$
where the inequality follows from the fact that $\phi$ is super-Gaussian. We may henceforth assume that $(Z_{t}(\theta))$ is upper-bounded by a nonnegative supermartingale for all $\theta\in\mathbb{B}^{d}$, enabling us to take $\Theta=\mathbb{B}^{d}$ in Proposition 2.2.
@@
$\int\log Z_{\tau}(\theta)\textnormal{d}\rho$ $=\lambda\langle\mu_{\rho},S_{\tau}\rangle-\psi(\lambda)\int\langle\theta,V_{\tau}\theta\rangle\textnormal{d}\rho$
$=\lambda\langle\mu_{\rho},S_{\tau}\rangle-\psi(\lambda)\left(\langle\mu_{\rho},V_{\tau}\mu_{\rho}\rangle+\text{Tr}(V_{\tau}\Sigma_{\rho})\right).$ (B.6)
For some $\beta>0$ to be determined later, we take $\rho$ to be the uniform distribution over the ellipsoid with mean $\mu_{\rho}=\beta V_{\tau}^{-1}S_{\tau}$ and shape $A_{\rho}=\psi^{-1}(\lambda)V_{\tau}^{-1}$. That is, $\rho$ is uniform over $\mathcal{E}_{\rho}:=\{\theta\in\mathbb{R}^{d}:(\theta-\mu_{\rho})^{t}A_{\rho}^{-1}(\theta-\mu_{\rho})\leq 1\}$ which we assume for now is a subset of $\mathbb{B}^{d}$. Note that similarly to the proof of Theorem 3.1, both $\mu_{\rho}$ and $A_{\rho}$ are functions of $\tau$ but we suppress this dependence for convenience. The distribution $\rho$ has covariance $\Sigma_{\rho}=(d+2)^{-1}A_{\rho}$ and
$\psi(\lambda)\text{Tr}(V_{\tau}\Sigma_{\rho})=\frac{d}{d+2}\leq 1.$
@@
$\int\log Z_{\tau}(\theta)\textnormal{d}\rho$ $=(\lambda\beta-\beta^{2}\psi(\lambda))\langle V_{\tau}^{-1}S_{\tau},S_{\tau}\rangle-\frac{d}{d+2}.$ (B.7)
@@
$(\lambda\beta-\beta^{2}\psi(\lambda))\|S_{\tau}\|_{V_{\tau}^{-1}}^{2}\leq D_{\textsc{kl}}(\rho\|\nu)+1+\log(1/\delta).$ (B.8)
@@
$D_{\textsc{kl}}(\rho\|\nu)=\frac{1}{2}\log\left(\frac{\det V_{\tau}}{\det U_{0}}\right).$ (B.9)
@@
It remains to pick $\beta$, which must be small enough such that $\mathcal{E}_{\rho}\subseteq\mathcal{E}_{\nu}$. That is, if $(\theta-\beta V_{\tau}^{-1}S_{t})^{t}A_{\rho}^{-1}(\theta-\beta V_{\tau}^{-1}S_{t})\leq 1$ we want to ensure that $\theta^{t}A_{\nu}^{-1}\theta\leq 1$. Equivalently, we may shift $\theta$ by $\beta V_{\tau}^{-1}S_{\tau}$ and show that if $\theta V_{\tau}\theta\leq\psi^{-1}(\lambda)$ then $(\theta+\beta V_{\tau}^{-1}S_{\tau})^{t}U_{0}(\theta+\beta V_{\tau}^{-1}S_{\tau})\leq\psi^{-1}(\lambda)$, where we’ve recalled the definition of $A_{\rho}$ and $A_{\nu}$. Put $y=y_{\tau}=V_{\tau}^{-1}S_{\tau}$. Suppose that $\theta^{t}V_{\tau}\theta\leq\psi^{-1}(\lambda)$. For any $\epsilon>0$,
@@
$\leq(1+\epsilon)\alpha_{\tau}\theta^{t}V_{\tau}\theta+\beta^{2}(1+\epsilon^{-1})y^{t}U_{0}y$ Rayleigh coef.
$\leq(1+\epsilon)\alpha_{\tau}\psi^{-1}(\lambda)+\beta^{2}(1+\epsilon^{-1})y^{t}U_{0}y$ $\theta^{t}V_{\tau}\theta\leq\psi^{-1}(\lambda)$
@@
so that $h(\epsilon^{*},\beta)=\beta^{2}\|y\|_{U_{0}}^{2}+2\beta\|y\|_{U_{0}}\sqrt{\alpha_{\tau}\psi^{-1}(\lambda)}+\alpha_{\tau}\psi^{-1}(\lambda)$. Setting $h(\epsilon^{*},\beta)=\psi^{-1}(\lambda)$ and solving for $\beta$ gives
$\beta^{*}=\frac{\sqrt{\alpha_{\tau}\psi^{-1}(\lambda)+1-\alpha_{\tau}}-\sqrt{\alpha_{\tau}\psi^{-1}(\lambda)}}{\|y\|_{U_{0}}}=\frac{g_{\tau}(\lambda)}{\|y\|_{U_{0}}}.$ (B.10)
Notice that $\|y\|^{2}_{U_{0}}=S_{\tau}^{t}V_{\tau}^{-1}U_{0}V_{\tau}^{-1}S_{\tau}\leq\|U_{0}\|\|S_{\tau}\|_{V_{\tau}^{-1}}^{2}$. Therefore, setting
$\beta^{**}=\frac{g_{\tau}(\lambda)}{\sqrt{\|U_{0}\|}\|S_{\tau}\|_{V_{\tau}^{-1}}}\leq\beta^{*},$
@@
$(\lambda\beta^{**}-(\beta^{**})^{2}\psi(\lambda))\|S_{\tau}\|_{V_{\tau}^{-1}}^{2}=\frac{\lambda g_{\tau}(\lambda)}{\sqrt{\|U_{0}\|}}\|S_{\tau}\|_{V_{\tau}^{-1}}-\frac{g_{\tau}^{2}(\lambda)\psi(\lambda)}{\|U_{0}\|},$
@@
Let $(S_{t},V_{t})$ be a sub-$\psi_{G,c}$ process for any $c>0$ and suppose that $V_{1}\succeq I_{d}$ and $V_{t}\succeq U_{0}$ for some positive-definite $U_{0}$. Let $\ell:\mathbb{N}_{0}\to\mathbb{R}_{>0}$ satisfy $\sum_{k\geq 0}\ell^{-1}(k)=1$. Fix $\delta\in(0,1)$ and take any $1<\eta<(e/\delta)^{2}$. Set
$\alpha_{\delta,\eta}:=1+\frac{\log(\eta)}{1+\log(1/\delta)-\log(\eta)/2}+\sup_{k\geq 0}\frac{\ell(k+2)}{\ell(k+1)}.$ (B.11)
Then, with probability $1-\delta$, for all stopping times $\tau$,
$\|S_{\tau}\|_{V_{\tau}^{-1}}\leq\frac{cD_{\tau,\eta}(U_{0},V_{\tau})+\sqrt{\frac{\alpha_{\delta,\eta}}{2}D_{\tau,\eta}(U_{0},V_{\tau})}+\max\left\{\frac{c+\sqrt{c^{2}+2\rho}}{2\rho},\sqrt{\frac{D_{\tau,\eta}(U_{0},V_{\tau})}{2}}\right\}}{H_{c}(U_{0},V_{\tau})},$ (B.12)
@@
$D_{\tau,\eta}(U_{0},V_{\tau})=\frac{1}{2}\log\left(\frac{\det V_{\tau}}{\det U_{0}}\right)+1+\log(\ell(\log_{\eta}(\det V_{\tau})))+\log(1/\delta),$ (B.13)
@@
$H_{c}(U_{0},V_{t})=0\vee\left(\sqrt{1-\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(V_{t})}}-\sqrt{\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(V_{t})}\frac{2}{(c+\sqrt{c^{2}+2c})}}\right).$ (B.14)
@@
Consider $\psi_{G,c}(\lambda)$ for $c>0$. We consider breaking intrinsic time into geometric epochs, $E_{k}=\{t:\eta^{k}\leq\det V_{t}<\eta^{k+1}\}$ for all $k\geq 0$ and some $\eta>1$. We assume that $V_{1}\succeq I$ so that $\cup_{k\geq 0}E_{k}$ is a bonafide partition of the sample space. Let $k(t)$ denote the unique $k\in\mathbb{N}$ such that $\eta^{k}\leq\det V_{t}\leq\eta^{k+1}$ and set
$D_{t}=\frac{1}{2}\log\left(\frac{\det V_{t}}{\det U_{0}}\right)+1+\log(1/\delta_{k(t)}).$ (B.15)
By definition, $k(t)\leq\log_{\eta}\det V_{t}$, so
@@
which is where the iterated logarithm term will come from in the final bound. In epoch $E_{k}$ we apply Theorem 4.1 with $\delta=\delta_{k}$ and $\lambda=\lambda_{k}$, for some $\delta_{k},\lambda_{k}$ to be chosen later. This gives that with probability $1-\delta_{k}$, for all $t\in E_{k}$,
$\|S_{t}\|_{V_{t}^{-1}}$ $\leq\frac{D_{t}}{g(\lambda_{k})\lambda_{k}}+\frac{g(\lambda_{k})\psi_{G,c}(\lambda_{k})}{\lambda_{k}}$
@@
$\leq\frac{1}{g(\lambda_{k})}\left(\frac{D_{t}}{\lambda_{k}}+\frac{\psi_{G,c}(\lambda_{k})}{\lambda_{k}}\right)=:W_{c}(t,\lambda_{k}),$
where we’ve used that $g(\lambda_{k})\leq 1$. Choose $\delta_{k}=\delta/\ell(k)$. Then,
$\mathbb{P}(\exists t\geq 1:\|S_{t}\|_{V_{t}^{-1}}\geq W_{c}(t,\lambda_{k(t)}))$ $=\mathbb{P}\bigg{(}\bigcup_{k\geq 0}\big{\{}\exists t\in E_{k}:\|S_{t}\|_{V_{t}^{-1}}\geq W_{c}(t,\lambda_{k})\big{\}}\bigg{)}$
$\leq\sum_{k\geq 0}\mathbb{P}\big{(}\exists t\in E_{k}:\|S_{t}\|_{V_{t}^{-1}}\geq W_{c}(t,\lambda_{k})\big{)}$
@@
$\lambda_{k}=\max\left\{\psi_{G,c}^{-1}(1/\rho),\frac{\sqrt{2B_{k}}}{1+c\sqrt{2B_{k}}}\right\},\text{~ where ~}\rho=\gamma_{\min}(U_{0}).$ (B.16)
Observe that $x/(1+cx)<1/c$ for all $x\geq 0$, so $\lambda_{k}<\lambda_{\max}$. Therefore, $\lambda_{k}\in[\psi_{G,c}^{-1}(1/\rho),\lambda_{\max})$ and is a legal choice in Theorem 4.1.
@@
$\alpha_{\delta,\eta}:=1+\frac{\log(\eta)}{1+\log(1/\delta)-\log(\eta)/2}+\sup_{k}\log(\ell(k+1)/\ell(k)).$ (B.17)
@@
where $s=\sup_{k}\log(\ell(k+1)/\ell(k))$. We claim that $\log(\eta)\leq uB_{k}$ for some constant $u$ and all $k$. Note, however, that we need only consider those $k$ such that $E_{k}$ actually occurs, i.e., the minimum $k$ we need to consider satisfies $\eta^{k}\leq\det V_{1}<\eta^{k+1}$. That is, we may assume that
$\frac{\eta^{k}}{\det U_{0}}>\frac{\det V_{1}}{\eta\det U_{0}}.$
@@
$\geq\frac{u}{2}\log\left(\frac{\det V_{1}}{\eta\det U_{0}}\right)+u+u\log(1/\delta)$
@@
where we’ve used that $V_{1}\succeq U_{0}$ by assumption. Therefore, to have $\log(\eta)\leq uB_{k}$ it suffices that $\log(\eta)\leq u\log(1/\eta)/2+u+u\log(1/\delta)$, i.e.,
@@
$B_{k+1}\leq B_{k}+\frac{u^{\circ}}{2}B_{k}+s\leq(1+u^{\circ}+s)B_{k}=\alpha_{\delta,\eta}B_{k}.$

### v2

@@ 7 Conclusion
The story does not end here. For one, as discussed in Section 4.3, we expect that our bounds are suboptimal, in the sense that they do not recover expected Bernstein-like behavior for $V_{t}=tI_{d}$. Can this issue be fixed? Second, can we remove the dependence on $H_{c,\tau}$ in Theorem 4.4? Is the function $g_{\psi}$ in Theorem 4.1 necessary, or is it simply an artifact of the proof technique? While $g_{\psi}$ is easy to compute, its presence makes it tougher to reason about these bounds analytically.
@@ Appendix A Method of Mixtures for sub-Gaussian Processes
Since $\int N_{t}(\theta)\textnormal{d}\nu(\theta)$ is a supermartingale by Fubini’s theorem, $M_{t}$ remains upper bounded by a nonnegative supermartingale. We may thus apply Ville’s inequality to obtain $P(M_{\tau}\geq 1/\delta)\leq\mathbb{E}[M_{1}]\delta\leq\delta$. In other words, with probability $1-\delta$, $\log M_{\tau}\leq\log(1/\delta)$, which translates to
@@ Appendix B Omitted Proofs
$\int\log Z^{\lambda}_{\tau}(\theta)\rho(\textnormal{d}\theta)$ $=\lambda\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}^{2}-\psi_{N}(\lambda)\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}V_{\tau}\widehat{V}_{\tau}^{-1}}^{2}-\psi_{N}(\lambda)\text{Tr}(V_{\tau}\Sigma_{\rho})$
@@
$=\frac{1}{2}\left\{\text{Tr}(\Sigma_{\nu}^{-1}\Sigma_{\rho})+\|S_{\tau}\|^{2}_{\widehat{V}_{\tau}^{-1}\Sigma_{\nu}^{-1}\widehat{V}_{\tau}^{-1}}-d+\log\left(\frac{\det\Sigma_{\nu}}{\det\Sigma_{\rho}}\right)\right\}.$
@@
First note that if $(S_{t},V_{t})$ is a sub-$\psi$ process, then so too is $(S_{t},V_{t}+U_{0})$ for any PSD $U_{0}$. Let $\widehat{V}_{t}=V_{t}+U_{0}$. For the remainder of the proof we will consider the process $(S_{t},\widehat{V}_{t})$. Set
$Z_{t}(\theta)=\exp\left\{\lambda\langle\theta,S_{t}\rangle-\psi(\lambda)\langle\theta,\widehat{V}_{t}\theta\rangle\right\}.$ (B.5)
For any $\theta\in\mathbb{S}^{d-1}$, the process $(Z_{t}(\theta))$ is upper bounded by some nonnegative supermartingale by assumption. We claim that the same holds for all $\theta\in\mathbb{B}^{d}$. This is shown in Chugg et al. [16, Lemma 2.10], but let us prove it for completeness. For any $\theta\in\mathbb{B}^{d}$, let $\phi=\theta/\|\theta\|$ and notice that
$\exp\{\lambda\langle\theta,S_{t}\rangle-\psi(\lambda)\langle\theta,\widehat{V}_{t}\theta\rangle\}$
$=\exp\{\lambda\|\theta\|\langle\phi,S_{t}\rangle-\psi(\lambda)\|\theta\|^{2}\langle\phi,\widehat{V}_{t}\phi\rangle\}$
$\leq\exp\{\lambda\|\theta\|\langle\phi,S_{t}\rangle-\psi(\|\theta\|\lambda)\langle\phi,\widehat{V}_{t}\phi\rangle\},$
where the inequality follows from the fact that $\psi$ is super-Gaussian. Since $\lambda\|\theta\|\leq\lambda$, $\psi(\|\theta\|\lambda)$ is well-defined and the final quantity defines a process which is upper bounded by a nonnegative supermartingale. We may henceforth assume that $(Z_{t}(\theta))$ is upper-bounded by a nonnegative supermartingale for all $\theta\in\mathbb{B}^{d}$, enabling us to take $\Theta=\mathbb{B}^{d}$ in Proposition 2.2.
@@
$\int\log Z_{\tau}(\theta)\textnormal{d}\rho$ $=\lambda\langle\mu_{\rho},S_{\tau}\rangle-\psi(\lambda)\int\langle\theta,\widehat{V}_{\tau}\theta\rangle\textnormal{d}\rho$
$=\lambda\langle\mu_{\rho},S_{\tau}\rangle-\psi(\lambda)\left(\langle\mu_{\rho},\widehat{V}_{\tau}\mu_{\rho}\rangle+\text{Tr}(\widehat{V}_{\tau}\Sigma_{\rho})\right).$ (B.6)
For some $\beta>0$ to be determined later, we take $\rho$ to be the uniform distribution over the ellipsoid with mean $\mu_{\rho}=\beta\widehat{V}_{\tau}^{-1}S_{\tau}$ and shape $A_{\rho}=\psi^{-1}(\lambda)\widehat{V}_{\tau}^{-1}$. That is, $\rho$ is uniform over $\mathcal{E}_{\rho}:=\{\theta\in\mathbb{R}^{d}:(\theta-\mu_{\rho})^{t}A_{\rho}^{-1}(\theta-\mu_{\rho})\leq 1\}$ which we assume for now is a subset of $\mathbb{B}^{d}$. Note that similarly to the proof of Theorem 3.1, both $\mu_{\rho}$ and $A_{\rho}$ are functions of $\tau$ but we suppress this dependence for convenience. The distribution $\rho$ has covariance $\Sigma_{\rho}=(d+2)^{-1}A_{\rho}$ and
$\psi(\lambda)\text{Tr}(\widehat{V}_{\tau}\Sigma_{\rho})=\frac{d}{d+2}\leq 1.$
@@
$\int\log Z_{\tau}(\theta)\textnormal{d}\rho$ $=(\lambda\beta-\beta^{2}\psi(\lambda))\langle\widehat{V}_{\tau}^{-1}S_{\tau},S_{\tau}\rangle-\frac{d}{d+2}.$ (B.7)
@@
$(\lambda\beta-\beta^{2}\psi(\lambda))\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}^{2}\leq D_{\textsc{kl}}(\rho\|\nu)+1+\log(1/\delta),$ (B.8)
@@
$D_{\textsc{kl}}(\rho\|\nu)=\frac{1}{2}\log\left(\frac{\det\widehat{V}_{\tau}}{\det U_{0}}\right).$ (B.9)
@@
It remains to pick $\beta$, which must be small enough such that $\mathcal{E}_{\rho}\subseteq\mathcal{E}_{\nu}$. That is, if $(\theta-\beta\widehat{V}_{\tau}^{-1}S_{t})^{t}A_{\rho}^{-1}(\theta-\beta\widehat{V}_{\tau}^{-1}S_{t})\leq 1$ we want to ensure that $\theta^{t}A_{\nu}^{-1}\theta\leq 1$. Equivalently, we may shift $\theta$ by $\beta V_{\tau}^{-1}S_{\tau}$ and show that if $\theta\widehat{V}_{\tau}\theta\leq\psi^{-1}(\lambda)$ then $(\theta+\beta\widehat{V}_{\tau}^{-1}S_{\tau})^{t}U_{0}(\theta+\beta\widehat{V}_{\tau}^{-1}S_{\tau})\leq\psi^{-1}(\lambda)$, where we’ve recalled the definition of $A_{\rho}$ and $A_{\nu}$. Put $y=y_{\tau}=\widehat{V}_{\tau}^{-1}S_{\tau}$ and suppose that $\theta^{t}\widehat{V}_{\tau}\theta\leq\psi^{-1}(\lambda)$. For any $\epsilon>0$,
@@
$\leq(1+\epsilon)\alpha_{\tau}\theta^{t}\widehat{V}_{\tau}\theta+\beta^{2}(1+\epsilon^{-1})y^{t}U_{0}y$ Rayleigh coef.
$\leq(1+\epsilon)\alpha_{\tau}\psi^{-1}(\lambda)+\beta^{2}(1+\epsilon^{-1})y^{t}U_{0}y$ $\theta^{t}\widehat{V}_{\tau}\theta\leq\psi^{-1}(\lambda)$
@@
so that $h(\epsilon^{*},\beta)=\beta^{2}\|y\|_{U_{0}}^{2}+2\beta\|y\|_{U_{0}}\sqrt{\alpha_{\tau}\psi^{-1}(\lambda)}+\alpha_{\tau}\psi^{-1}(\lambda)$. Setting $h(\epsilon^{*},\beta)=\psi^{-1}(\lambda)$ and solving for $\beta$ gives (after taking the positive root of the quadratic equation)
$\beta^{*}:=\frac{\sqrt{\psi^{-1}(\lambda)}-\sqrt{\alpha_{\tau}\psi^{-1}(\lambda)}}{\|y\|_{U_{0}}}=\frac{g_{\tau}(\lambda)}{\|y\|_{U_{0}}}.$ (B.10)
Notice that
$\|y\|^{2}_{U_{0}}$ $=S_{\tau}^{t}\widehat{V}_{\tau}^{-1}U_{0}\widehat{V}_{\tau}^{-1}S_{\tau}=(S_{\tau}^{t}\widehat{V}_{\tau}^{-1/2})(\widehat{V}_{\tau}^{-1/2}U_{0}\widehat{V}_{\tau}^{-1/2})(\widehat{V}_{\tau}^{-1/2}S_{\tau})$
$\leq\|\widehat{V}_{\tau}^{-1/2}U_{0}\widehat{V}_{\tau}^{-1/2}\|S_{\tau}^{t}\widehat{V}_{\tau}^{-1}S_{\tau}\leq\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}^{2},$
where the final inequality follows since $U_{0}\preceq\widehat{V}_{\tau}$ so $\widehat{V}_{\tau}^{-1/2}U_{0}\widehat{V}_{\tau}^{-1/2}\preceq I$ and
$\|\widehat{V}_{\tau}^{-1/2}U_{0}\widehat{V}_{\tau}^{-1/2}\|\leq 1.$
Therefore, setting
$\beta^{**}:=\frac{g_{\tau}(\lambda)}{\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}}\leq\beta^{*},$
@@
$(\lambda\beta^{**}-(\beta^{**})^{2}\psi(\lambda))\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}^{2}=\lambda g_{\tau}(\lambda)\|S_{\tau}\|_{\widehat{V}_{\tau}^{-1}}-g_{\tau}^{2}(\lambda)\psi(\lambda),$
@@
Let $(S_{t},V_{t})$ be a sub-$\psi_{G,c}$ process for any $c>0$ and suppose that $V_{1}+U_{0}\succeq I_{d}$. Let $U_{0}$ be positive definite and let $\ell:\mathbb{N}_{0}\to\mathbb{R}_{>0}$ satisfy $\sum_{k\geq 0}\ell^{-1}(k)=1$. Fix $\delta\in(0,1/e]$ and take any $1<\eta<(e/\delta)^{2}$. Set $\upsilon=\gamma_{\min}(U_{0})$,
$\alpha_{\delta,\eta}:=1+\frac{\log(\eta)}{2+2\log(1/\delta)-\log(\eta)}+\sup_{k\geq 0}\frac{\ell(k+2)}{\ell(k+1)},$ (B.11)
and $\xi(x)=2/(c+\sqrt{c^{2}+2x})$. Then, with probability $1-\delta$, for all stopping times $\tau$,
$\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}\leq\frac{cD_{\tau,\eta}(U_{0},V_{\tau})+\sqrt{\frac{\alpha_{\delta,\eta}}{2}D_{\tau,\eta}(U_{0},V_{\tau})}+\xi(c)\max\left\{\frac{1}{\upsilon\xi(\upsilon)},\sqrt{\frac{D_{\tau,\eta}(U_{0},V_{\tau})}{2}}\right\}}{H_{c}(U_{0},V_{\tau})},$ (B.12)
@@
$D_{\tau,\eta}(U_{0},V_{\tau})=\frac{1}{2}\log\left(\frac{\det(V_{\tau}+U_{0})}{\det U_{0}}\right)+1+\log(\ell(\log_{\eta}(\det(V_{\tau}+U_{0}))))+\log(1/\delta),$ (B.13)
@@
$H_{c}(U_{0},V_{t})=0\vee\left(\sqrt{\frac{\sqrt{c^{2}+2c}-c}{\sqrt{\alpha_{\delta,\eta}}+c}}-\sqrt{\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(\widehat{V}_{t})}\xi(c)}\right)$ (B.14)
@@
Let $\widehat{V}_{t}=V_{t}+U_{0}$. Consider $\psi_{G,c}(\lambda)$ for $c>0$. We consider breaking intrinsic time into geometric epochs, $E_{k}=\{t:\eta^{k}\leq\det\widehat{V}_{t}<\eta^{k+1}\}$ for all $k\geq 0$ and some $\eta>1$. Recall we are assuming that $U_{0}\succeq I_{d}$, implying that $\cup_{k\geq 0}E_{k}$ is a bonafide partition of the sample space. Let $k(t)$ denote the unique $k\in\mathbb{N}$ such that $\eta^{k}\leq\det\widehat{V}_{t}\leq\eta^{k+1}$ and set
$D_{t}=\frac{1}{2}\log\left(\frac{\det\widehat{V}_{t}}{\det U_{0}}\right)+1+\log(1/\delta_{k(t)}).$ (B.15)
Note that $k(t)$ is a random variable. By definition, $k(t)\leq\log_{\eta}\det\widehat{V}_{t}$, so
@@
which is where the iterated logarithm term will come from in the final bound. In epoch $E_{k}$ our strategy is to apply Theorem 4.1 with $\delta=\delta_{k}$ and $\lambda=\lambda_{k}$, for some $\delta_{k},\lambda_{k}$ to be chosen later. For $g=g_{\psi,t}$, this gives that with probability $1-\delta_{k}$, for all $t\in E_{k}$,
$\|S_{t}\|_{\widehat{V}_{t}^{-1}}$ $\leq\frac{D_{t}}{g(\lambda_{k})\lambda_{k}}+\frac{g(\lambda_{k})\psi_{G,c}(\lambda_{k})}{\lambda_{k}}$
@@
$\leq\frac{1}{g(\lambda_{k})}\left(\frac{D_{t}}{\lambda_{k}}+\frac{\xi(c)\psi_{G,c}(\lambda_{k})}{\lambda_{k}}\right)=:W_{c}(t,\lambda_{k}),$
where we’ve used that $g(\lambda_{k})\leq\sqrt{\psi_{G,c}^{-1}(\lambda_{k})}\leq\sqrt{\xi(c)}$ since $\psi_{G,c}^{-1}(u)$ is increasing in $u$ and $\lambda_{k}<\lambda_{\max}$. Choose $\delta_{k}=\delta/\ell(k)$. Then,
$\mathbb{P}(\exists t\geq 1:\|S_{t}\|_{V_{t}^{-1}}\geq W_{c}(t,\lambda_{k(t)}))$ $=\mathbb{P}\bigg(\bigcup_{k\geq 0}\big\{\exists t\in E_{k}:\|S_{t}\|_{\widehat{V}_{t}^{-1}}\geq W_{c}(t,\lambda_{k})\big\}\bigg)$
$\leq\sum_{k\geq 0}\mathbb{P}\big(\exists t\in E_{k}:\|S_{t}\|_{\widehat{V}_{t}^{-1}}\geq W_{c}(t,\lambda_{k})\big)$
@@
$\lambda_{k}=\max\left\{\psi_{G,c}^{-1}(1/\upsilon),\frac{\sqrt{2B_{k}}}{1+c\sqrt{2B_{k}}}\right\},\text{~ where ~}\upsilon=\gamma_{\min}(U_{0}).$ (B.16)
Observe that $x/(1+cx)<1/c$ for all $x\geq 0$, so $\lambda_{k}<\lambda_{\max}$. Therefore, $\lambda_{k}\in[\psi_{G,c}^{-1}(1/\upsilon),\lambda_{\max})$ and is a legal choice in Theorem 4.1.
@@
$\alpha_{\delta,\eta}:=1+\frac{\log(\eta)}{2+2\log(1/\delta)-\log(\eta)}+\sup_{k}\log(\ell(k+1)/\ell(k)).$ (B.17)
@@
where $s=\sup_{k}\log(\ell(k+1)/\ell(k))$. We claim that $\log(\eta)\leq uB_{k}$ for some constant $u$ and all $k$. Note, however, that we need only consider those $k$ such that $E_{k}$ actually occurs, i.e., the minimum $k$ we need to consider satisfies $\eta^{k}\leq\det\widehat{V}_{1}<\eta^{k+1}$. That is, we may assume that
$\frac{\eta^{k}}{\det U_{0}}>\frac{\det\widehat{V}_{1}}{\eta\det U_{0}}.$
@@
$\geq\frac{u}{2}\log\left(\frac{\det\widehat{V}_{1}}{\eta\det U_{0}}\right)+u+u\log(1/\delta)$
@@
where we’ve used that $\widehat{V}_{1}\succeq U_{0}$. Therefore, to have $\log(\eta)\leq uB_{k}$ it suffices that $\log(\eta)\leq u\log(1/\eta)/2+u+u\log(1/\delta)$, i.e.,
@@
$B_{k+1}\leq B_{k}+\frac{u^{\circ}}{2}B_{k}+s=\alpha_{\delta,\eta}B_{k}.$